1) A student reads a barometer on the wall and the pressure reading is 776 mmHg. This is equivalent to how many kilo pascals of pressure? *


2) The air in an automobile tire exerts 200.0 kPa when the temperature is 15.0 degrees Celsius. Assuming constant volume, what will be the pressure the air exerts when the temperature rises to 30.0 degrees Celsius? *


3) In order to condense oxygen gas into liquid oxygen,one research facility decided to cool pure oxygen to gas to 60.0 degrees Kelvin and pressure it to 2.50 atm. What would these conditions be equivalent to in degrees Celsius and millimeter of mercury? *


1
Expert's answer
2020-06-29T07:52:26-0400

1) 1 mmHg = 133.322 pascals

  1 kilo pascales  = 1000 pascals

776 mmHg = (776×133.322) / 1000 kilo pascales  

                = 103.45 kilo pascales  (Answer)     

2) 15.0 degrees Celsius = 273.15+15 Kelvin

                                        =288.15 Kelvin

30.0 degrees Celsius = 303.15 Kelvin

At constant volume,  P1/T1 = P2/T2

P1 = 200.0 kPa, T1 = 288 Kelvin,  T2 =303 Kelvin


P2 = (P1× T2) / T1

   = (200.0× 303.15) / 288.15 kPa

  = 210.4 kPa

The air exerts 210.4 kPa  (Answer)

3) 0 Kelvin = -273.15 degrees Celsius

  60 Kelvin = -213.15 degrees Celsius (Answer)

1 atm = 760 millimeter of mercury

 2.50 atm = (760×2.50) millimeter of mercury

              = 1900 millimeter of mercury (Answer)